OpenStudy (anonymous):

6 11 -- = -- x+2 x Solve for x? <3

myininaya (myininaya):

6x=11(x+2) 6x=11x+22 6x-11x=22 -5x=22 x=-22/5

OpenStudy (gw2011):

You can cross multiply and you get: 11(x+2)=6x 11x+22=6x 5x=-22 x=-22/5=-4 2/5=-4.4

OpenStudy (anonymous):

okay thank you, i figured out i just wasnt multiplying the 11 and the 2. xD thanks so much!